OpenStudy (anonymous):

@ranga @Becki A new car has a sticker price of $22,450, while the invoice price paid was $19,450. What is the percentage markup? A. 15.42% B. 11.36% C. 13.36% D. 10%

OpenStudy (anonymous):

Subtract the invoice price from the sticker price. Then divide that number by the invoice price. That's your percentage.

OpenStudy (anonymous):

b

OpenStudy (anonymous):

That isn't what I got. 22450-19450=x x/19450=percentage markup

OpenStudy (anonymous):

idk its confusing me

OpenStudy (anonymous):

OK. So first, subtract the dealer price from the sticker price.

OpenStudy (anonymous):

i got 6.4

OpenStudy (anonymous):

22450-19450=

OpenStudy (anonymous):

3,000

OpenStudy (anonymous):

Now divide that by 19450, the price the dealer paid (invoice).

OpenStudy (anonymous):

a

OpenStudy (anonymous):

yes. :)